12. In question 8, assume that, beginning from the initial equilibrium position (investment equal to 100, government expenditure equal to 75, and net taxes fixed at 100), there was an autonomous fall in consumption and an increase in saving such that the consumption function shifted from C = 25 + 0.8YD to C = 5 + 0.8YD a. Find the change in equilibrium income resulting from this autonomous increase in saving. b. Calculate the level of saving before and after the shift in the consumption and, therefore, the saving function. How do you explain this result?
1
Expert's answer
2019-10-08T09:20:39-0400

a. The change in equilibrium income resulting from this autonomous increase in saving is: (25 - 5)/0.8 = 25.

b. The formula for saving is S = Y - C = I + G.

The level of saving before the shift and after the shift in the consumption is the same, because S = 100 + 75 = 175 in both cases.